OpenStudy (anonymous):

What is the x-coordinate of the vertex of the parabola whose equation is y = 3x2 + 9x?

OpenStudy (anonymous):

the x-coordinate of the vertex of y = ax^2 + bx + c is given by: x = -b/(2a).

OpenStudy (anonymous):

y = 3x^2 + 9x = 3x^2 + 9x + 0

OpenStudy (anonymous):

a = 3, b = 9, and c = 0

OpenStudy (anonymous):

the x-coordinate of the vertex is: x = -b/(2a) = -9/[2(3)] = -3/2 = -1 1/2

OpenStudy (anonymous):

-3 -1 - this are the answer choices @EllenJaz17

OpenStudy (anonymous):

What's the last option? It's just a negative sign

OpenStudy (anonymous):

no its -2 over 3 -2/3

OpenStudy (anonymous):

and so is the second answer is -1 and 1 over 2 somethuing like this -1 1/2

OpenStudy (anonymous):

Its -3/2

OpenStudy (anonymous):

y = 3x^2 + 9x y = 3(x^2 + 3x + 9/4) - 27/4 y = 3(x + 3/2)^2 - 27/4 From this, we can see a low point and vertex when: x = -3/2 because that's when the first term is smallest. So, x = -3/2

OpenStudy (anonymous):

thanks:)